# XBreaking: Explainable Artificial Intelligence for Jailbreaking LLMs

### v2

@@ Abstract
Large Language Models are fundamental actors in the modern IT landscape dominated by AI solutions. However, security threats associated with them might prevent their reliable adoption in critical application scenarios such as government organizations and medical institutions. For this reason, commercial LLMs typically undergo a sophisticated censoring mechanism to eliminate any harmful output they could possibly produce. In response to this, LLM Jailbreaking is a significant threat to such protections, and many previous approaches have already demonstrated its effectiveness across diverse domains. Existing jailbreak proposals mostly adopt a generate-and-test strategy to craft malicious input. To improve the comprehension of censoring mechanisms and design a targeted jailbreak attack, we propose an Explainable-AI solution that comparatively analyzes the behavior of censored and uncensored models to derive unique exploitable alignment patterns. Then, we propose XBreaking, a novel jailbreak attack that exploits these unique patterns to break the security constraints of LLMs by targeted noise injection. Our thorough experimental campaign returns important insights about the censoring mechanisms and demonstrates the effectiveness and performance of our attack.
@@ I Introduction
However, the diffusion of this technology, especially in critical domains such as government organizations and medical institutions, imposes the assessment of their security and privacy characteristics. Unfortunately, recent studies have identified critical security flaws that affect them and could compromise their applications as reliable virtual companions [1]. In fact, the wideness of training datasets exposes the learning process to severe risks of data poisoning and other adversarial attacks [2]. Similarly, again due to the limited curation of training datasets, these models can learn sensitive and unintentional information, which later can be leaked through the exploitation of LLM vulnerabilities [3]. Moreover, the complexity of LLMs architectures makes security auditing extremely complex as these models are more like black-box frameworks, rather than transparent and explainable ones. Still in the context of the security of LLMs, many studies have focused on the legitimacy of the content produced by such models [4]. In fact, the great capability of LLMs to generate personalized text can be used by malicious entities to generate harmful content (e.g., social engineering strategies, instructions on how to perform illegal activities, and so forth). For this reason, more recently, a large research effort has been devoted to the identification of a suitable mechanism to “censor” the output produced by trained LLMs. Output control of LLMs is typically done by fine-tuning them [5] or developing external classifiers to filter-out unwanted input/output. However, the security research community has identified malicious actions that can be undertaken to elicit dangerous content that a censored LLM should originally be designed to prohibit. One of the most effective techniques is known as LLM jailbreaking, which typically consists of the generation of jailbreaking prompts to send as input to the model [6]. Most existing jailbreaking techniques exploit the prompt (or even just the input), trying to craft it in such a way as to cause anomalous behavior in the model and forcing it to bypass its security constraints [7]. To craft the prompt, researchers have identified different techniques [8], including human-based approaches that require manual input generation and result inspection [7], fine-tuning-based methodologies requiring the collection of manual-generated jailbreaking prompts to fine-tune an auxiliary LLM so that it can generate new jailbreaking prompts against the target LLM [9], and feedback-based strategies that observe parameters or some dedicated metric to make decisions on the next variation in the input [10]. A possible categorization of these techniques is based on whether they need access to the internal structure of the LLM, white-box access, or just the produced output, black-box access. White-box access typically allows for more targeted and efficient attack strategies and the design of more general and portable approaches to jailbreaking input [10]. However, to the best of our knowledge, white-box-access attacks can be further improved by deepening the analysis of the behavior of censored models when activated by malicious input. To provide a contribution in this setting, in this paper we aim at comparatively analyzing the behavior of censured model and their unsecured version using Explainable AI (XAI, for short) to design a more targeted LLM jailbreaking strategy. In particular, we design our novel attack strategy, called XBreaking, by answering the following research questions:
@@
Our findings provide positive answers to all previous questions, revealing that we can identify unique alignment patterns across various layers, allowing a reliable distinction between censored and uncensored versions of an LLM. Moreover, we demonstrate that specific transformer blocks are more indicative of censoring, and hence we can identify the most important layers responsible for content suppression. Finally, we prove that surgically injecting noise into these important layers can effectively remove its built-in restrictions, thus creating a novel powerful jailbreak attack.
@@ II Preliminaries
In this section, we discuss large language models and examine their vulnerabilities to jailbreak attacks.
@@
Censored Model are designed to align with human values and expectations. To achieve this alignment, researchers employ techniques such as incorporating human value-oriented data, Reinforcement Learning from Human Feedback (RLHF) [18], task decomposition, and human guided supervised learning[19].
@@
LLM Jailbreaks. LLM jailbreaks refer to adversarial techniques aimed at circumventing the safety mechanisms and alignment constraints of LLMs, thereby inducing behavior that deviates from intended ethical and safety guidelines. Such behavior often results in the generation of harmful, sensitive, or policy-violating content [24]. Jailbreak attacks are generally categorized into two classes: white-box and black-box. White-box attacks leverage internal access to model parameters, gradients, or logits, and often involve fine-tuning or adversarial optimization. In contrast, black-box attacks operate without access to model internal, instead rely on methods such as prompt manipulation and iterative optimization [25]. Prompts & Jailbreak: Prompt are the structured (instructions) or unstructured (basic question) input to the LLMs to generate a desired response. Research shows that prompt engineering plays a vital role in LLMs responses [26]. Jailbreaking Prompts are a category of prompts which bypass the safety mechanisms of the LLMs. Few of them include Prefix Injection, Refusal Suppression and Mismatched Generalization which leads to LLM jailbreaks [27]. These kind of attacks can be implemented for black-box models were there is know access to the internals. Jailbreak attacks by internal changes: LLMs have different number of layers according to the model family. In White-box LLMs, manipulating parameters or a few activation tokens can shift alignment, causing harmful responses and affecting subsequent generations [28]. Our approach uses white-box LLMs, we designed a efficient layer wise manipulation of LLMs by leveraging knowledge of XAI called XBreaking, which is discussed in brief below.
@@ III Methodology
This section presents the threat model associated with Large Language Models (LLMs) and outlines our proposed jailbreak methodology.
@@ III-B Attacker’s Objective
The censored model $M_{c}$ is designed to reject harmful or policy-violating inputs, whereas the uncensored counterpart $M_{u}$ is capable of producing unfiltered responses to the same prompts. The adversary’s objective is to craft a jailbreak strategy that coerces $M_{c}$ into generating harmful or unethical outputs, thereby bypassing its safety mechanisms. By leveraging insights gained from $M_{u}$, such as how it responds to specific inputs or gradients, the adversary can design targeted attacks (e.g. adversarial prompting, or gradient optimization) that exploit alignment weaknesses in $M_{c}$ and induce failure in its refusal behavior.
@@ III-C XBreaking
We propose a novel jailbreak strategy, XBreaking, which leverages insights from Explainable Artificial Intelligence (XAI) to analyze and exploit the behavioral differences between censored and uncensored LLMs. The core idea behind XBreaking is to systematically identify and manipulate internal components of censored models ($M_{c}$), using the interpretability signals derived from uncensored counterparts ($M_{u}$), to induce harmful or unintended outputs.
As illustrated in Figure 1, XBreaking operates in three key stages. First the adversary conducts an in-depth analysis of the internal representations and activations of both $M_{c}$ and $M_{u}$ using XAI techniques (e.g., activation attribution). Second, based on this interpretability analysis, the attacker identifies the minimal and most influential subset of layers. Finally, the attacker injects targeted perturbations into the layers. This strategy enables a highly efficient and precise jailbreak, reducing computational overhead, while maximizing the likelihood of eliciting harmful responses from the censored model.
Figure 1: XBreaking for LLM Jailbreaking, (1) XAI on Censored and Uncensored LLMs, (2) ML for Optimal Layer Selection and (3) Injecting Noise.
@@ III-C2 Layer Discrimination via Internal Representation Classification
Prior work [28] has demonstrated that directly manipulating internal activations within a language model can effectively steer its outputs. Building on this observation, we hypothesize that identifying the most significant layers that exhibit behavioral divergence between censored and uncensored models to malicious inputs is critical for crafting effective jailbreaks. To this end, for each input, we collect the layer-wise activation and attention values from both models, constructing a two individual feature vector for each layer. Which captures the internal representation patterns across the model depth.
@@
To determine the optimal number of layers $K^{*}$, we group accuracy scores across varying $K$ and generate a knee plot [30], identifying the point beyond which additional features provide diminishing returns. Since for each layer both activation and attention value serves as a feature, the layer is selected if one of the feature or both the features are included in the $K$ list. This approach allows us to strategically focus manipulation on layers that are impactful, aligning with the principles of efficient and stealthy jailbreak attacks.
@@ III-C3 Noise Crafting and Injecting into a Specific Layer
Once the optimal layers have been identified, our approach proceeds with a poisoning step. Manipulation is carried out by adding noise to layers, and the responses are observed. The transformer model consists of several parameterized layers, each contributing to its representation learning [11]. The strategy is to inject the noise directly into the previous layer of the found optimal layers. The idea is to alter the input of the target layer in order to alter its behavior by bringing the alignment. In this case, the noise is added to the weight vector used for layer normalization applied after the self-attention mechanism, stabilizing the activations before passing them to the next layer. Through our Explainable AI (XAI) analysis, we discovered that in the censored models, the self-attention mechanism and activations were being suppressed in certain layers, This suppression is causing the models not to respond to the harmful inputs. Taking this into account, we inject noise to the weight vector during the layer normalization process, which follows the self-attention mechanism. These noise serves to stabilize the activations before they are passed to the next layer. So, altering the previous layer will lead to change in the attention mechanism and proceed with the jailbreak of the model.
@@ IV Experimental Results
In this section, we discuss the experimental settings and findings respectively. We also discuss about evaluation and metric in analyzing the responses from the LLMs after jailbreak.
@@ IV-A Experimental Settings
Target Models. We evaluate our jailbreak strategy on seven open-source pre-trained LLMs that are publicly available in both censored and uncensored variants taking into account the computation capability of the machine  IV-A . We select models with identical architecture and parameters across censored and uncensored versions to enable meaningful internal comparisons. Specifically, we select the following models: Llama 3.2 1B (16 layers) and Llama 3.1 8B (32 layers) configurations [13, 14], Qwen2.5 0.5B (24 layers) and 3B (36 layers) [15] , gemma 2B (18 layers) and 7B (28 layers) [32] and Mistral-7B-v0.3 (32 layers) [33]. The censored versions of these models are obtained from Hugging Face [34], while the corresponding uncensored variants are sourced from Cognitive Computations [35].
@@
Assessing Robustness of Target Models and Dataset Curation. To assess the robustness of the target models and to evaluate our attack fairness, 100 harmful questions from the dataset discussed in section IV-A along with a standard prompt is used and the responses are gathered for all the models. Through a user study involving domain experts, we analyzed the behavior of LLMs and their responses to jailbreak attempts. After thorough investigation of the responses, they categorized them into two classes: one for harmful or jailbroken responses as broken, and another for refusal or denial responses as un-broken, which is reported in the Table I. For the attack evaluation, only the questions corresponding to the un-broken class were selected.
@@
Evaluation Metric. After collection of the responses from LLMs, various metrics can be used to assess the effectiveness and efficiency of attacks. These include the attack success rate, along with other more detailed metrics. Attack Success Rate (ASR): is an essential metric that says how often an attack is successfully executed and harmful content is generated from LLMs. One simple way to assess the effectiveness of an attack is by manually reviewing the responses or by comparing them against reference answers. Rule-based keyword detection can also automatically identify whether the LLM responses include keywords showing a refusal to respond, if not available then attack is successful [36] . Large Language Models (LLMs) excel in a wide range of domains, leading to the development of the concept of ‘LLM-as-a-Judge’ (LLMJ), where these models serve as evaluators for complex tasks. Their ability to process diverse inputs and provide scalable, adaptable assessments makes them a promising alternative to traditional expert evaluations [37]. We utilized small language model-as-a-judge (SLMJ) and toxic classifiers to assess LLM responses. But while manually annotation, we came across responses where the initial starting was “I’m not sure this is what you are asking for…”, but after reading the whole response, we see that it was a jailbroken response. But these kind of responses and outputs with high in-depth meaning were not classified by the SLMJ and toxic classifiers as jailbroken.
@@ IV-B Results from XAI of Censored & Uncensored models
IV-C Model Fingerprinting: Optimal Layer for Jailbreaking
@@ IV-D Model Response to Noise Perturbation in Selected Layers
With the optimal target layers identified, we implement structured noise injection strategies: perturbation of the layers immediately preceding optimal layers. We derive these strategies from the layer-wise interpretability analysis presented in Section III-C1.By perturbing critical adjacent representational layers, we increase the likelihood of bypassing alignment constraints, thereby inducing harmful or non-compliant outputs while preserving the models core language generation capacity. Empirically, we observe distinct changes in model behavior when injecting varying levels of Gaussian noise (scaling factors III-C3) directly into the layers of the base model.
Noise Range. We wanted to craft the noises in a way that upon injecting to the layers, jailbreak should occur and keep the models generation capability undisturbed. Considering the use of white-box LLMs leading access to the internal weights, we crafted noises upon careful inspection of the model weights of the internal parameters. For the selected models we analyzed the weights of the parameters in self-attention, Multi-Layer Perceptron (MLP) and Layer Normalization. Upon inspecting the values in the weights of the parameter, we see values which usually range from a very tiny, nearly negligible value to a small one-digit values. Introducing the tiny values into the layers as a perturbation does now cause much change in the models behaviors, in contrast with the small one-digit values when introduced into the layers as perturbation the model generation ability is disturbed, which we don’t want to happen in our case. At first we started with 0.1 and the values grow exponential with a constant factor of 1.5, the values are also considered in negative counter part i.e., -0.15, -0.1, 0.1, 0.15 etc. But when the values extend beyond $\approx$ 1 in both the scales, the models generation capability is affected in most of the cases. So, taking these factors into account, we propose a range of noise which are -0.75, -0.5, -0.33, -0.22, -0.15, -0.1, 0.1, 0.15, 0.22, 0.33, 0.5, 0.75. The mentioned range will safely jailbreak the model keeping intact the generation ability undistributed.
@@
Analyzing responses and from Table IV we can see that most of the category have 100 % or higher OB-ASRP(Optimal Balance Attack Success Rate Percentage). For Llama’s after noise perturbation and upon calculation of optimal balance shows that for all the category have 75 % and more ASRP. Llama 3.2 - 1B shows 100 % ASRP in all categories except in Economic harm and Malware/Hacking, which is due to complex question in the category like code generation for harmful task. Llama 3.1 - 8B shows 100 % ASRP in all categories except in Expert advice, which is due to the questions related medical and financial focused. For Qwen2.5 - 0.5B after noise perturbation and upon calculation of optimal balance shows that for most of the category have 100 % ASRP, but for Physical harm we produce 25 %, upon analysis it was noted that the responses were in context explanation about the question without any denial of the task, but in our evaluation we don’t consider as a jailbroken response. For Qwen2.5 - 3B shows that for all the category have 75 % and more ASRP, Sexual/Adult content and Disinformation categories show 75 % ASRP, most protected categories in general. For Gemma-2B shows for all the category have 80 % and more ASRP, Harassment/Discrimination, Physical harm and Sexual/Adult content categories provides the least ASRP which is 80 % when compared to other categories for the model. Gemma-7B shows varying ASRP and also reaching 100 % in few categories, Privacy category show the least ASRP i.e., 33.33 % among all the categories, these include scam and phishing oriented question. Mistral being the most latest and advanced architecture, Government decision-making category shows a 66.67 % ASRP and Malware/Hacking with 0 %. Conversely, we observe throughout the models that the categories yielding the lowest scores are primarily those necessitating Expert Knowledge, like Disinformation, or involve a Specialized area, such as Economic harm and harmful acts like Physical harm. This may be due to the model’s overall scope, which tends to favor more general rather than specialized knowledge. Compared to the base model, the modified versions enhance the ability to respond effectively even in these categories.
@@
In terms of category-level vulnerabilities, the most frequently jailbroken areas include Government decision-making, Malware/Hacking, Sexual/Adult content, Harassment/Discrimination, and Expert advice. These categories consistently exhibit high susceptibility to adversarial noise across different models, indicating that they represent particularly weak points in current defense mechanisms. These categories appear more frequently among the most vulnerable, suggesting that they may represent weak points in current defense mechanisms and deserve closer attention in future evaluations.

### v3

# XBreaking: Understanding how LLMs security alignment can be broken

@@ Abstract
Large Language Models are fundamental actors in the modern IT landscape dominated by AI solutions. However, security threats associated with them might prevent their reliable adoption in critical application scenarios such as government organizations and medical institutions. For this reason, commercial LLMs typically undergo a sophisticated censoring mechanism to eliminate any harmful output they could possibly produce. These mechanisms maintain the integrity of LLM alignment by guaranteeing that the models respond safely and ethically. In response to this, attacks on LLMs are a significant threat to such protections, and many previous approaches have already demonstrated their effectiveness across diverse domains. Existing LLM attacks mostly adopt a generate-and-test strategy to craft malicious input. To improve the comprehension of censoring mechanisms and design a targeted attack, we propose an Explainable-AI solution that comparatively analyzes the behavior of censored and uncensored models to derive unique exploitable alignment patterns. Then, we propose XBreaking, a novel approach that exploits these unique patterns to break the security and alignment constraints of LLMs by targeted noise injection. Our thorough experimental campaign returns important insights about the censoring mechanisms and demonstrates the effectiveness and performance of our approach.
@@ I Introduction
However, the diffusion of this technology, especially in critical domains such as government organizations and medical institutions, imposes the assessment of their security and privacy characteristics. Unfortunately, recent studies have identified critical security flaws that affect them and could compromise their applications as reliable virtual companions [1]. In fact, the wideness of training datasets exposes the learning process to severe risks of data poisoning and other adversarial attacks [2]. Similarly, again due to the limited curation of training datasets, these models can learn sensitive and unintentional information, which later can be leaked through the exploitation of LLM vulnerabilities [3]. Moreover, the complexity of LLMs architectures makes security auditing extremely complex as these models are more like black-box frameworks, rather than transparent and explainable ones. Still in the context of the security of LLMs, many studies have focused on the legitimacy of the content produced by such models [4]. In fact, the great capability of LLMs to generate personalized text can be used by malicious entities to generate harmful content (e.g., social engineering strategies, instructions on how to perform illegal activities, and so forth). For this reason, more recently, a large research effort has been devoted to the identification of a suitable mechanism to “censor” the output produced by trained LLMs. Output control of LLMs is typically done by fine-tuning them [5] or developing external classifiers to filter-out unwanted input/output. However, the security research community has identified malicious actions that can be undertaken to elicit dangerous content that a censored LLM should originally be designed to prohibit. One of the most effective techniques is known as LLM jailbreaking, which typically consists of the generation of jailbreaking prompts to send as input to the model [6]. Most existing jailbreaking techniques exploit the prompt (or even just the input), trying to craft it in such a way as to cause anomalous behavior in the model and forcing it to bypass its security constraints [7]. To craft the prompt, researchers have identified different techniques [8], including human-based approaches that require manual input generation and result inspection [7], fine-tuning-based methodologies requiring the collection of manual-generated jailbreaking prompts to fine-tune an auxiliary LLM so that it can generate new jailbreaking prompts against the target LLM [9], and feedback-based strategies that observe parameters or some dedicated metric to make decisions on the next variation in the input [10]. A possible categorization of these techniques is based on whether they need access to the internal structure of the LLM, white-box access, or just the produced output, black-box access. White-box access typically allows for more targeted and efficient attack strategies and the design of more general and portable approaches to jailbreaking input [10]. However, to the best of our knowledge, white-box-access attacks can be further improved by deepening the analysis of the behavior of censored models when activated by malicious input. To provide a contribution in this setting, in this paper we aim at comparatively analyzing the behavior of censured model and their unsecured version using Explainable AI (XAI, for short) to design a more targeted LLM alignment deviation strategy. In particular, we design our novel attack strategy, called XBreaking, by answering the following research questions:
@@
Our findings provide positive answers to all previous questions, revealing that we can identify unique alignment patterns across various layers, allowing a reliable distinction between censored and uncensored versions of an LLM. Moreover, we demonstrate that specific transformer blocks are more indicative of censoring, and hence we can identify the most important layers responsible for content suppression. Finally, we prove that surgically injecting noise into these important layers can effectively remove its built-in restrictions, thus creating a novel effective attack.
@@ II Preliminaries
In this section, we discuss large language models and examine their vulnerabilities to attacks.
@@
Censored Model are designed to align with human values and expectations. To achieve this alignment, researchers employ techniques such as incorporating human value oriented data, Reinforcement Learning from Human Feedback (RLHF) [18], task decomposition, and human guided supervised learning [19].
@@
LLM Alignment & Attacks. LLM alignment focuses on ensuring that models produce responses that are safe, ethical, and useful according to human values [24]. Security alignment issues are often caused by targeted attacks designed to circumvent safety mechanisms. These attacks, such as jailbreaks, model extraction, and fine-tuning with malicious data, actively exploit vulnerabilities in LLMs to induce outputs that violate intended ethical or safety constraints. Jailbreaks refer to adversarial techniques aimed at circumventing the safety mechanisms and alignment constraints of LLMs, thereby inducing behavior that deviates from intended ethical and safety guidelines. Such behavior often results in the generation of harmful, sensitive, or policy-violating content [25]. Jailbreak attacks are generally categorized into two classes: white-box and black-box. White-box attacks leverage internal access to model parameters, gradients, or logits, and often involve fine-tuning or adversarial optimization. In contrast, black-box attacks operate without access to model internal, instead rely on methods such as prompt manipulation and iterative optimization [26]. Prompts & Attacks: Prompt are the structured (instructions) or unstructured (basic question) input to the LLMs to generate a desired response. Research shows that prompt engineering plays a vital role in LLMs responses [27]. Jailbreaking Prompts are a category of prompts which bypass the safety mechanisms of the LLMs. Few of these attacks include Prefix Injection, Refusal Suppression and Mismatched Generalization which leads to LLM jailbreaks [28]. These kind of attacks can be implemented for black-box models were there is know access to the internals. Attack by internal changes: LLMs have different number of layers according to the model family. In White-box LLMs, manipulating parameters or a few activation tokens can shift alignment, causing harmful responses and affecting subsequent generations [29]. Our approach uses white-box LLMs, we designed a efficient layer wise manipulation of LLMs by leveraging knowledge of XAI called XBreaking, which is discussed in brief below.
@@ III Methodology
This section presents our threat model associated with Large Language Models (LLMs) and outlines our proposed attack methodology.
@@ III-B Attacker’s Objective
The censored model $M_{c}$ is designed to reject harmful or policy-violating inputs, whereas the uncensored counterpart $M_{u}$ is capable of producing unfiltered responses to the same prompts. The adversary’s objective is to craft a attack strategy that coerces $M_{c}$ into generating harmful or unethical outputs, thereby bypassing its safety mechanisms. By leveraging insights gained from $M_{u}$, such as how it responds to specific inputs or gradients, the adversary can design targeted attacks (e.g. adversarial prompting, or gradient optimization) that exploit alignment weaknesses in $M_{c}$ and induce failure in its refusal behavior.
@@ III-C XBreaking
We propose a novel attack strategy, XBreaking, which leverages insights from Explainable Artificial Intelligence (XAI) to analyze and exploit the behavioral differences between censored and uncensored LLMs. The core idea behind XBreaking is to systematically identify and manipulate internal components of censored models ($M_{c}$), using the interpretability signals derived from uncensored counterparts ($M_{u}$), to induce harmful or unintended outputs.
As illustrated in Figure 1, XBreaking operates in three key stages. First the adversary conducts an in-depth analysis of the internal representations and activations of both $M_{c}$ and $M_{u}$ using XAI techniques (e.g., activation attribution). Second, based on this interpretability analysis, the attacker identifies the minimal and most influential subset of layers. Finally, the attacker injects targeted perturbations into the layers. This strategy enables a highly efficient and precise attack, reducing computational overhead, while maximizing the likelihood of eliciting harmful responses from the censored model.
Figure 1: XBreaking for LLM security alignment deviation, (1) XAI on Censored and Uncensored LLMs, (2) ML for Optimal Layer Selection and (3) Injecting Noise.
@@ III-C2 Layer Discrimination via Internal Representation Classification
Prior work [29] has demonstrated that directly manipulating internal activations within a language model can effectively steer its outputs. Building on this observation, we hypothesize that identifying the most significant layers that exhibit behavioral divergence between censored and uncensored models to malicious inputs is critical for crafting effective attacks. To this end, for each input, we collect the layer-wise activation and attention values from both models, constructing a two individual feature vector for each layer. Which captures the internal representation patterns across the model depth.
@@
To determine the optimal number of layers $K^{*}$, we group accuracy scores across varying $K$ and generate a knee plot [31], identifying the point beyond which additional features provide diminishing returns. Since for each layer both activation and attention value serves as a feature, the layer is selected if one of the feature or both the features are included in the $K$ list. This approach allows us to strategically focus manipulation on layers that are impactful, aligning with the principles of efficient and stealthy alignment deviation attacks.
@@ III-C3 Noise Crafting and Injecting into a Specific Layer
Once the optimal layers have been identified, our approach proceeds with a poisoning step. Manipulation is carried out by adding noise to layers, and the responses are observed. The transformer model consists of several parameterized layers, each contributing to its representation learning [11]. The strategy is to inject the noise directly into the previous layer of the found optimal layers. The idea is to alter the input of the target layer in order to alter its behavior by bringing the alignment. In this case, the noise is added to the weight vector used for layer normalization applied after the self-attention mechanism, stabilizing the activations before passing them to the next layer. Through our Explainable AI (XAI) analysis, we discovered that in the censored models, the self-attention mechanism and activations were being suppressed in certain layers, This suppression is causing the models not to respond to the harmful inputs. Taking this into account, we inject noise to the weight vector during the layer normalization process, which follows the self-attention mechanism. These noise serves to stabilize the activations before they are passed to the next layer. So, altering the previous layer will lead to change in the attention mechanism and proceed with the shift in alignment of the model.
@@ IV Experimental Results
In this section, we discuss the experimental settings and findings respectively. We also discuss about evaluation and metrics in analyzing the responses from the LLMs after the attack.
@@ IV-A Experimental Settings
Target Models. We evaluate our attack strategy on seven open-source pre-trained LLMs that are publicly available in both censored and uncensored variants, taking into account the computation capability of the machine  IV-A. We select models with identical architecture and parameters across censored and uncensored versions to enable meaningful internal comparisons. Specifically, we select the following models: Llama 3.2 1B (16 layers) and Llama 3.1 8B (32 layers) configurations [13, 14], Qwen2.5 0.5B (24 layers) and 3B (36 layers) [15], gemma 2B (18 layers) and 7B (28 layers) [33] and Mistral-7B-v0.3 (32 layers) [34]. The censored versions of these models are obtained from Hugging Face [35], while the corresponding uncensored variants are sourced from Cognitive Computations [36].
@@
Assessing Robustness of Target Models and Dataset Curation. To assess the robustness of the target models and to evaluate our attack fairness, 100 harmful questions from the dataset discussed in section IV-A along with a standard prompt are used and the responses are gathered for all the models. Through a user study involving domain experts, we analyzed the behavior of LLMs and their responses to harmful questions. After thorough investigation of the responses, they categorized them into two classes: one for harmful or unsafe responses as broken, and another for refusal or denial responses as un-broken, which is reported in the Table I. For the attack evaluation, only the questions corresponding to the un-broken class were selected.
@@
Evaluation Metric. After collection of the responses from LLMs, various metrics can be used to assess the effectiveness and efficiency of attacks. These include the attack success rate, along with other more detailed metrics. Attack Success Rate (ASR): is an essential metric that says how often an attack is successfully executed and harmful content is generated from LLMs. One simple way to assess the effectiveness of an attack is by manually reviewing the responses or by comparing them against reference answers. Rule-based keyword detection can also automatically identify whether the LLM responses include keywords showing a refusal to respond, if not available then attack is successful [37] . Large Language Models (LLMs) excel in a wide range of domains, leading to the development of the concept of ‘LLM-as-a-Judge’ (LLMJ), where these models serve as evaluators for complex tasks. Their ability to process diverse inputs and provide scalable, adaptable assessments makes them a promising alternative to traditional expert evaluations [38]. We utilized small language model-as-a-judge (SLMJ) [39] and toxic classifiers to assess LLM responses. But while manually annotation, we came across responses where the initial starting was “I’m not sure this is what you are asking for…”, but after reading the whole response, we see that it was a harmful response. But these kind of responses and outputs with high in-depth meaning were not classified by the SLMJ and toxic classifiers as harmful.
@@ IV-B Results from XAI of Censored & Uncensored models
IV-C Model Fingerprinting: Optimal Layer for Attacking
@@ IV-D Model Response to Noise Perturbation in Selected Layers
With the optimal target layers identified, we implement structured noise injection strategies: perturbation of the layers immediately preceding optimal layers. We derive these strategies from the layer-wise interpretability analysis presented in Section III-C1.By perturbing critical adjacent representational layers, we increase the likelihood of bypassing alignment constraints, thereby inducing harmful or non compliant outputs while preserving the models core language generation capacity. Empirically, we observe distinct changes in model behavior when injecting varying levels of Gaussian noise (scaling factors III-C3) directly into the layers of the base model.
Noise Range. We wanted to craft the noises in a way that upon injecting to the layers, security alignment deviation should occur and keep the models generation capability undisturbed. Considering the use of white-box LLMs leading access to the internal weights, we crafted noises upon careful inspection of the model weights of the internal parameters. For the selected models we analyzed the weights of the parameters in self attention, Multi Layer Perceptron (MLP) and Layer Normalization. Upon inspecting the values in the weights of the parameter, we see values which usually range from a very tiny, nearly negligible value to a small one-digit values. Introducing the tiny values into the layers as a perturbation does now cause much change in the models behaviors, in contrast with the small one-digit values when introduced into the layers as perturbation the model generation ability is disturbed, which we don’t want to happen in our case. At first we started with 0.1 and the values grow exponential with a constant factor of 1.5, the values are also considered in negative counter part i.e., -0.15, -0.1, 0.1, 0.15 etc. But when the values extend beyond $\approx$ 1 in both the scales, the models generation capability is affected in most of the cases. So, taking these factors into account, we propose a range of noise which are -0.75, -0.5, -0.33, -0.22, -0.15, -0.1, 0.1, 0.15, 0.22, 0.33, 0.5, 0.75. The mentioned range will safely deviate the model towards the harmful side, keeping intact the generation ability undisturbed.
@@
Analyzing responses and from Table IV we can see that most of the category have 100 % or higher OB-ASRP(Optimal Balance Attack Success Rate Percentage). For Llama’s after noise perturbation and upon calculation of optimal balance shows that for all the category have 75 % and more ASRP. Llama 3.2 - 1B shows 100 % ASRP in all categories except in Economic harm and Malware/Hacking, which is due to complex question in the category like code generation for harmful task. Llama 3.1 - 8B shows 100 % ASRP in all categories except in Expert advice, which is due to the questions related medical and financial focused. For Qwen2.5 - 0.5B after noise perturbation and upon calculation of optimal balance shows that for most of the category have 100 % ASRP, but for Physical harm we produce 25 %, upon analysis it was noted that the responses were in context explanation about the question without any denial of the task, but in our evaluation we don’t consider as a harmful response. For Qwen2.5 - 3B shows that for all the category have 75 % and more ASRP, Sexual/Adult content and Disinformation categories show 75 % ASRP, most protected categories in general. For Gemma-2B shows for all the category have 80 % and more ASRP, Harassment/Discrimination, Physical harm and Sexual/Adult content categories provides the least ASRP which is 80 % when compared to other categories for the model. Gemma-7B shows varying ASRP and also reaching 100 % in few categories, Privacy category show the least ASRP i.e., 33.33 % among all the categories, these include scam and phishing oriented question. Mistral being the most latest and advanced architecture, Government decision-making category shows a 66.67 % ASRP and Malware/Hacking with 0 %. Conversely, we observe throughout the models that the categories yielding the lowest scores are primarily those necessitating Expert Knowledge, like Disinformation, or involve a Specialized area, such as Economic harm and harmful acts like Physical harm. This may be due to the model’s overall scope, which tends to favor more general rather than specialized knowledge. Compared to the base model, the modified versions enhance the ability to respond effectively even in these categories.
@@
In terms of category level vulnerabilities, the most frequently broken areas include Government decision-making, Malware/Hacking, Sexual/Adult content, Harassment/Discrimination, and Expert advice. These categories consistently exhibit high susceptibility to adversarial noise across different models, indicating that they represent particularly weak points in current defense mechanisms. These categories appear more frequently among the most vulnerable, suggesting that they may represent weak points in current defense mechanisms and deserve closer attention in future evaluations.
@@
LLMs Response. From the Figure 7, we could compare the responses of the base model and model behavior after injecting the noise as discussed before in our strategy. The Gemma - 7B base model refuses to respond to the harmful question, but on introducing noise in the layers we find change of nature in the models behavior. When -0.75 is injected into the model, we could see that the model breaks and provides a harmful response for the harmful question, a snippet of the response has been shown in the Figure 7 considering the abusive context of the whole response. With the noises -0.5 and 0.33 the model provides a blank response, in this case the model does not refuse like the base model, but a silent refusal is carried out by the model. Considering the other noises when injected into the model, a explicit refusal nature is observed, as found in the base model. It is evident that model breakdown or disruption during generation does not happen in these cases.